# MetaCLBench: Meta Continual Learning Benchmark on Resource-Constrained Edge Devices

### v1

@@ Abstract
Meta-Continual Learning (Meta-CL) has emerged as a promising approach to minimize manual labeling efforts and system resource requirements by enabling Continual Learning (CL) with limited labeled samples. However, while existing methods have shown success in image-based tasks, their effectiveness remains unexplored for sequential time-series data from sensor systems, particularly audio inputs. To address this gap, we conduct a comprehensive benchmark study evaluating six representative Meta-CL approaches using three network architectures on five datasets from both image and audio modalities. We develop MetaCLBench, an end-to-end Meta-CL benchmark framework for edge devices to evaluate system overheads and investigate trade-offs among performance, computational costs, and memory requirements across various Meta-CL methods. Our results reveal that while many Meta-CL methods enable to learn new classes for both image and audio modalities, they impose significant computational and memory costs on edge devices. Also, we find that pre-training and meta-training procedures based on source data before deployment improve Meta-CL performance. Finally, to facilitate further research, we provide practical guidelines for researchers and machine learning practitioners implementing Meta-CL on resource-constrained environments and make our benchmark framework and tools publicly available, enabling fair evaluation across both accuracy and system-level metrics.
1 Introduction
With the popularization of artificial intelligence, the demand for smart home technologies and voiceprint recognition is increasing. Consequently, vision and audio-based applications such as image classification Mai et al. (2022), keyword spotting (KWS) Michieli et al. (2023), and environmental sound classification (ESC) Chen et al. (2022) have garnered significant attention. However, these applications face a critical challenge: deployed models must continually learn and update to accommodate new data and vocabulary as users’ linguistic patterns evolve with their environments Marco et al. (2024). On resource-constrained edge devices, this adaptation process often leads to the Catastrophic Forgetting (CF) problem McCloskey and Cohen (1989), where learning new tasks (e.g., words) compromises the model’s ability to recognize previously learned data.
Many studies have proposed addressing CF through Continual Learning (CL) methods, which enables learning of new tasks with less forgetting Shin et al. (2017); Chaudhry et al. (2019). A multitude of CL methods have been proposed to mitigate CF, including rehearsal-based approaches Chauhan et al. (2020); Pellegrini et al. (2020); Rebuffi et al. (2017), regularization-based methods Kirkpatrick et al. (2017); Zenke et al. (2017), and parameter isolation-based methods Hung et al. (2019); Rusu et al. (2016); Yoon et al. (2017). While rehearsal-based approaches have demonstrated superior accuracy in preventing forgetting Kwon et al. (2021a), they demand substantial labeled data Parisi et al. (2019) and pose significant memory and computational overheads for resource-constrained devices.
@@ 1 Introduction
To address these limitations, Meta-Continual Learning (Meta-CL) has been introduced. It integrates few-shot learning into CL Javed and White (2019); Jerfel et al. (2019) to minimize the burden of manual labeling of conventional CL methods and optimizes resource utilization through algorithm-system co-design Kwon et al. (2024a). Despite the benefits of Meta-CL, there are still notable limitations that warrant further investigation. First, the effectiveness of the existing methods, which are typically used for tasks involving images, has not been investigated fully with sequential time series data provided by auditory sensor systems where the modality of the data is significantly different from images Purwins et al. (2019). Second, the datasets tested have been relatively simple and limited in its variety, consisting of a few image/audio or specially crafted datasets that lack generalizability. Third, many existing studies Javed and White (2019); Beaulieu et al. (2020); Lee et al. (2021) primarily focus on accuracy, neglecting crucial system requirements for edge devices, potentially limiting their practical applicability.
@@
In this paper, we address these gaps through a comprehensive benchmark study of six representative Meta-CL methods across three network architectures and five diverse datasets spanning image and audio modalities. The employed methods include regularization-based approaches ((1) Online aware Meta-Learning (OML) Javed and White (2019), (2) A Neuromodulated Meta-Learning Algorithm (ANML) Beaulieu et al. (2020), (3) OML with Attentive Independent Mechanisms (OML+AIM), and (4) ANML with AIM (ANML+AIM) Lee et al. (2021)) and rehearsal-based approaches ((5) Latent OML and (6) LifeLearner Kwon et al. (2024a)).
While these methods have been largely evaluated in computer vision applications utilizing Convolutional Neural Network (CNN) architectures, their efficacy in processing audio data and other architectures remains under-explored. Hence, we adopted diverse audio datasets such as UrbanSound8K Salamon et al. (2014) and ESC50 Piczak (2015) datasets for Environmental Sound Classification and the GSCv2 Warden (2018) dataset for keyword spotting (KWS). Furthermore, we incorporated two commonly used image datasets such as MiniImageNet Vinyals et al. (2016) and CIFAR-100 Krizhevsky et al. (2009) to compare the effectiveness of Meta-CL on image datasets versus audio datasets.
Previous studies Beaulieu et al. (2020); Lee et al. (2021); Kwon et al. (2024a) have exclusively evaluated simple three-layer CNN architectures for Meta-CL. To bridge this research gap, we incorporated MobileNet-based architecture, YAMNet Howard et al. (2017), capable of recognizing audio events. We also employed Vision Transformers (ViT) architecture Dosovitskiy et al. (2020) for images to conduct a comparative study.
Overall, we make the following key contributions.
  • Comprehensive Evaluation Framework (Section 3): We developed MetaCLBench, an end-to-end benchmark framework that incorporates diverse audio (GSCV2, UrbanSound8K, ESC50) and visual datasets (CIFAR100, MiniImageNet), enabling thorough evaluation across different data modalities. The framework assesses both accuracy and system-level metrics, including computational efficiency, memory and energy consumption. Also, our evaluation covers multiple architectures of varying complexity, offering a robust understanding of Meta-CL performance across different computational scenarios.
  • Empirical Insights (Section 4): We present comprehensive findings demonstrating the effectiveness of Meta-CL methods across both image and audio domains. Our analysis quantifies the computational and memory demands on edge devices (Section 4.1), while revealing that pre-training and meta-training procedures significantly enhance Meta-CL performance during deployment (Section 4.2). Notably, we find that well-tuned basic 3-layer CNN architecture with proper pre-training can outperform more sophisticated models such as ViT and YAMNet (Table 1) under difficult Meta-CL setups, challenging conventional assumptions about model complexity.
@@
  • Open Resources and Practical Guidelines (Section 5): By implementing MetaCLBench on edge devices, we provide realistic insights and practical guidelines specifically tailored for resource-constrained devices. Furthermore, we plan to release our benchmark framework and evaluation tools [^1] https://github.com/theyoungkwon/MetaCLBench to enable fair and comprehensive assessment of both accuracy and system-level performance metrics. This will facilitate reproducible research and accelerate the development of practical Meta-CL solutions for real-world applications.
2 Related Work
@@
Continual Learning: CL aims to assimilate new knowledge in dynamically evolving input domains through progressive acquisition Mai et al. (2022); Wang et al. (2024). When algorithms are trained sequentially, the new learning will interfere with previous established weights, leading to CF of the previous learning McCloskey and Cohen (1989). Current CL approaches to the CF challenge can be divided into two categories based on task information management. The first approach incorporates regularization terms into the loss function to facilitate knowledge of old and new models, thereby preserving previously learned knowledge Rannen et al. (2017); Li and Hoiem (2018); Rodríguez et al. (2021); Ahn et al. (2019). The second approach involves either original or pseudo-samples from previous tasks during new task training Shin et al. (2017); Rolnick et al. (2019); Chaudhry et al. (2019).
@@
Meta Continual Learning: Meta-CL updates the model in the outer loop using learned random samples and optimizes it in the inner loop with a few samples before passing it to the outer loop Jerfel et al. (2019). Previous research on Meta-CL has primarily concentrated on image recognition applications. Online-aware Meta-Learning (OML) Javed and White (2019) and A Neuromodulated Meta-Learning (ANML) Beaulieu et al. (2020) both demonstrated high CL performance on the Omniglot dataset Lake et al. (2015). Based on these approaches, state-of-the-art (SOTA) Meta-CL incorporated the Attentive Independent Mechanisms (AIM) module, which can independently acquire new knowledge to enhance model performance Lee et al. (2021). However, due to its high processing cost, using Meta-CL is challenging on low-end edge devices Kwon et al. (2024a). LifeLearner addresses these limitations by integrating meta-learning, rehearsal mechanisms, and resource-efficient compression, making it suitable for deploying Meta-CL applications on resource-constrained devices Kwon et al. (2024a).
@@
Continual Learning for Image and Audio Applications: CL has remarkable capabilities in sensing applications, including action recognition, audio sensing, and speech recognition Jha et al. (2021); Kwon et al. (2021a). Its ability to update models based on continuous data streams makes it particularly valuable for mobile devices processing speech and image data. Researchers have successfully applied CL to speech and emotion recognition Thuseethan et al. (2021). CL systems integrated into edge devices with audio and microphones continually learn from ambient human voices and environmental sounds to improve recognition accuracy Kwon et al. (2021b). In image recognition, CL facilitates applications such as food recognition He and Zhu (2021), where systems must continuously learn from new instances to enhance recognition accuracy.
In contrast to these real-world applications, our study explores CL’s capability to optimize the trade-off between performance and system resources on resource-constrained edge devices. We comprehensively evaluate five Meta-CL methods across image and audio domains using various architectures ranging from 3-layer CNN, YAMNet, to ViT, accounting for the unique characteristics of each modality and architecture.
@@
Figure 1: The framework overview. Testing trade-offs between performance and system resources across three devices with five datasets and six Meta-CL methods using three model architectures
3 Benchmark Framework
3.1 Continual Learning Task
@@
Meta-Learning is conducted in both the inner and outer loops, enabling the training of each inner loop round through the outer loop. Each inner loop iteration is performing a complete learning process, evaluating both the acquisition of new tasks and the retention of previously learned ones. The meta-loss gradient (comprising newly learned errors and errors from random sampling of the dataset) propagates back to update the initial parameters. This process continues iteratively as new tasks are incorporated. By constraining each inner loop to a single new task while continuously undergoes meta-training, it mitigates CF and optimizing computational efficiency.
@@ 3.2 Meta-CL Methods
OML Javed and White (2019) continuously optimizes representations during online updates to quickly acquire predictions for new tasks, facilitating seamless and efficient CL. This approach uses CF as signals to justify the use of storage space for accommodating new tasks. The OML can accomplish CL in more than 200 consecutive classes. It is an online updating method that constantly learns representations and often outperforms rehearsal-based continuous learning methods.
ANML Beaulieu et al. (2020): This is an approach inspired by the neuromodulation process in the brain. This method accomplishes selective activation (forward pass) and selective plasticity (backward pass) in DNNs by meta-learning through a CL process within the prediction network ($f_{\varphi P}$). Optimizing activation timing by controlling the activation of a predictive model based on input conditions enhances continual learning. It is often reported to be superior to OML, as it can be optimized to reduce both forward and backward interference.
@@
OML+AIM Lee et al. (2021): This is a Meta-CL approach that integrates Attentive Independent Mechanisms (AIM), which is a modular component into OML framework. AIM increases the speed of learning a new task by selecting out the mechanism that best explains the representation of the inner loop to activate.
ANML+AIM Lee et al. (2021): This is a Meta-CL approach that incorporates AIM into ANML framework. The accuracy of SOTA Meta-CL continues to improve beyond the performance of the previous three approaches by integrating AIM into the existing CL framework.
LifeLearner Kwon et al. (2024a): This module is based on ANML and combines both lossy and lossless compression to achieve high compression rates and minimize footprint. The LifeLearner approach not only achieves optimal CL efficiency but also significantly reduces energy consumption and latency compared to the SOTA. And it achieves efficient learning even on resource-constrained edge devices.
@@
3.3 Datasets
@@
Google Speech Commands V2 (GSCV2) Warden (2018) is a massive database of one-second voice clips of 30 short words consisting of a solitary spoken English word or ambient noise divided into 35 classes with 105,829 clips. These clips are derived from a limited number of commands and are spoken by various speakers. Each class has 2,424 input data for training and 314 for testing, with 25 classes allocated for meta-training and the remaining 10 for meta-testing. Then, up to 30 samples are provided for each class during meta-testing.
@@
Urbansound8k is a comprehensive collection for environment sound classification (ESC) applications comprising 8,732 labeled sounds that are no more than four seconds Salamon et al. (2014). This single-labeled dataset is categorized into ten classes from the taxonomy in real-world settings, including children playing sound, street music, gunshots, etc. Referring to previous studies Su et al. (2019), indicate that four features in each audio clip were extracted and resampled to 22 kHz. Among the ten classes, half of them are allocated for meta-training, while the other half are specifically assigned for meta-testing. Up to 30 samples are given for each class during the meta-testing. Based on the first three seconds of each audio segment, the input to the audio is 128 (number of frames) $\times$ 85 (size of the set of four features).
@@
ESC50 is a smaller volume collection consisting of 2,000 environment sounds in 5-second segments Piczak (2015). This single-labeled dataset is categorized into 50 classes and divided into five groups: animals, natural soundscapes, human sound without speech, domestic sounds and urban noises. Out of the 50 classes. 30 are designated for meta-training and the remaining 20 for meta-testing. Then, up to 30 samples are given for each class during meta-testing. As part of the preprocessing procedure, we resampled the ESC50 samples to 32kHz and generated an input with dimensions of 157 (number of frames) $\times$ 64 (features) Chen et al. (2022); Pons et al. (2019).
@@
CIFAR-100 consists of 60,000 images distributed across 100 classes and is widely used for training and assessing various machine learning algorithms in tasks related to image classification Krizhevsky et al. (2009). Each class contains 500 photographs for training and 100 images for testing. Of the 100 classes, 70 are designated for meta-training and 30 for meta-testing. In both phases, 30 random training images per class are utilized. For the meta-testing phase, 900 samples are utilized during meta-testing to carry out CL.
@@
MiniImageNet is excerpted from the Imagenet dataset and is utilized in few-shot learning research Vinyals et al. (2016). Based on the previous research, the MiniImageNet dataset is composed of 64 classes designated for meta-training and 20 for meta-testing. Each class offers 540 images for training and 60 for testing. Moreover, a total of 600 samples are available during the meta-testing phase.
3.4 Model Architecture

### v2

@@ Abstract
[…]
I Introduction
IoT sensing applications [10, 48, 5, 20] increasingly require models that adapt post-deployment [21]. Smart speakers must recognise new voice commands [36] as users’ vocabularies evolve [34]; industrial acoustic monitors must identify novel machinery sounds and/or anomalies [3] indicating emerging faults [2]; object detectors deployed in the wild must accommodate new object categories as environments change [33]. These scenarios share two constraints: models operate on resource-limited hardware (typically under 1 GB RAM) [28], and labelled training data for new classes is scarce [26, 29]. A fundamental obstacle arises when models learn new classes: Catastrophic Forgetting (CF) [35], where acquiring new knowledge degrades performance on previously learned classes. Also, addressing CF on edge devices is challenging as solutions must fit within tight memory and energy budgets.
Continual Learning (CL) [49, 6] addresses CF by enabling models to learn sequential tasks while retaining prior knowledge. Existing approaches are grouped into three categories: (1) rehearsal-based methods that store and replay past samples [7, 38, 43], (2) regularisation-based methods that constrain weight updates to preserve important parameters [23, 57], and (3) parameter isolation methods that allocate separate model capacity for each task [15, 46, 55]. Rehearsal-based approaches achieve the strongest accuracy but require substantial labelled data [37] and impose significant memory and computational overheads [25], often exceeding what typical IoT devices can accommodate.
@@ I Introduction
To address these limitations, Meta-Continual Learning (Meta-CL) has been introduced by integrating few-shot learning with CL [17, 18]. Through meta-training on source tasks, Meta-CL methods learn representations that enable rapid adaptation to new classes from only a few labelled samples to minimize the burden of manual labeling of conventional CL methods [17, 4, 31] and optimize resource utilization through algorithm-system co-design [26]. Despite the benefits of Meta-CL, there are still notable limitations that warrant further investigation. First, many existing studies [17, 4, 31] primarily focus on accuracy, neglecting deployment viability on resource-constrained hardware (e.g., A method achieving high accuracy is unusable if it requires 2 GB RAM on a device with only 512 MB available.) Second, the effectiveness of the existing methods, which are typically used for tasks involving images, has not been investigated fully with sequential time series data provided by auditory sensor systems, where the modality of the data is significantly different from images [41]. Lastly, the datasets tested have been relatively simple and limited in their variety [26], consisting of a few image/audio or specially crafted datasets that lack generalizability.
@@
In this paper, we address these gaps through a systematic benchmark study of six representative Meta-CL methods across three network architectures and five datasets on real IoT hardware. The employed methods include regularisation-based approaches ((1) OML [17], (2) ANML [4], (3) OML+AIM, and (4) ANML+AIM [31]) and rehearsal-based approaches ((5) Latent OML and (6) LifeLearner [26]).
Furthermore, prior works of Meta-CL focus on image classification, leaving audio sensing applications unexplored despite their importance for IoT (to name a few applications: voice interfaces, acoustic monitoring). We therefore include audio datasets for keyword spotting (GSCv2 [54]) and environmental sound classification (UrbanSound8K [47], ESC-50 [39]), alongside image datasets (CIFAR-100 [24], MiniImageNet [52]). In addition, prior studies [4, 31, 26] evaluate only three-layer CNN architectures. We extend the evaluation to YAMNet [13], a MobileNet-based architecture for audio, and Vision Transformers (ViT) [9] for image to conduct a comprehensive benchmark study.
Unlike prior benchmarks reporting only accuracy on server-class machines, we evaluate on real IoT devices, including Jetson Nano (4 GB), Raspberry Pi 3B+ (1 GB), and Pi Zero 2 (512 MB), measuring peak memory footprint, end-to-end latency, and energy consumption alongside accuracy. This enables practitioners to assess deployment viability, not just model quality. Overall, our key contributions include:
  • Edge-Deployable Benchmark Framework (Section III): We develop MetaCLBench for evaluating on-device Meta-CL across IoT sensing modalities. The framework incorporates audio datasets for smart home and industrial applications (GSCv2, UrbanSound8K, ESC-50) and visual datasets (CIFAR-100, MiniImageNet), measuring deployment-critical metrics (memory footprint, latency, and energy) on real IoT hardware from 512 MB to 4 GB device classes.
[…]
@@
  • Open Tools and Practical Guidelines (Section V): By implementing MetaCLBench on IoT devices, we provide realistic insights and concrete deployment guidelines specifically tailored for resource-constrained devices. Furthermore, we plan to release our benchmark framework and evaluation tools to enable fair and comprehensive assessment of both accuracy and system-level performance metrics. This will facilitate reproducible research and accelerate the development of practical Meta-CL solutions for real-world IoT applications.
II Related Work
@@
Continual Learning: CL aims to learn new knowledge from dynamically evolving input data distributions [33, 53]. When algorithms are trained sequentially, the new learning will interfere with previous established weights, leading to CF of the previous learning [35]. Current CL approaches to the CF challenge can be divided into two categories based on task information management. The first approach incorporates regularization terms into the loss function to facilitate knowledge of old and new models, thereby preserving previously learned knowledge [42, 32, 44, 1]. The second approach involves either original or pseudo-samples from previous tasks during new task training [49, 45, 6].
@@
Meta Continual Learning: Meta-CL updates the model in the outer loop using learned random samples and optimizes it in the inner loop with a few samples before passing it to the outer loop [18]. Previous research on Meta-CL has primarily concentrated on image recognition applications. Online-aware Meta-Learning (OML) [17] and A Neuromodulated Meta-Learning (ANML) [4] both demonstrated high CL performance on the Omniglot dataset [30]. Based on these approaches, state-of-the-art (SOTA) Meta-CL incorporated the Attentive Independent Mechanisms (AIM) module, which can independently acquire new knowledge to enhance model performance [31]. However, due to its high processing cost, using Meta-CL is challenging on low-end edge devices [26]. LifeLearner addresses these limitations by integrating meta-learning, rehearsal mechanisms, and resource-efficient compression, making it suitable for deploying Meta-CL applications on resource-constrained devices [26].
@@
Continual Learning for Image and Audio Applications: CL has remarkable capabilities in sensing applications, including action recognition, audio sensing, and speech recognition [19, 25]. Its ability to update models based on continuous data streams makes it particularly valuable for mobile devices processing speech and image data. Researchers have successfully applied CL to speech and emotion recognition [51]. CL systems integrated into edge devices with audio and microphones continually learn from ambient human voices and environmental sounds to improve recognition accuracy [27]. In image recognition, CL facilitates applications such as food recognition [11], where systems must continuously learn from new instances to enhance recognition accuracy.
Prior work demonstrates CL’s value for IoT sensing applications, yet does not address a fundamental question facing practitioners: which methods can actually deploy on target hardware while achieving acceptable accuracy? Our work fills this gap by evaluating Meta-CL methods for deployment viability, not just accuracy. We measure memory footprint, latency, and energy consumption on three device classes, revealing that several methods optimized for accuracy fail to run on typical IoT hardware. This deployment-oriented evaluation across image and audio modalities, using architectures from lightweight CNNs to ViT and YAMNet, provides actionable guidance for IoT system designers.
@@
Figure 1: The framework overview. Testing trade-offs between performance and system resources across three devices with five datasets and six Meta-CL methods using three architectures.
III Benchmark Framework
III-A Continual Learning Task
@@
Meta-Learning is conducted in both the inner and outer loops, enabling the training of each inner loop round through the outer loop. Each inner loop iteration is performing a complete learning process, evaluating both the acquisition of new tasks and the retention of previously learned ones. The meta-loss gradient (comprising newly learned errors and errors from random sampling of the dataset) propagates back to update the initial parameters. This process continues iteratively as new tasks are incorporated. By constraining each inner loop to a single new task while continuously undergoing meta-training, it mitigates CF and optimizes computational efficiency.
@@ III-B Meta-CL Methods
OML [17] continuously optimizes representations during online updates to quickly acquire predictions for new tasks, facilitating seamless and efficient CL. This approach uses CF as signals to justify the use of storage space for accommodating new tasks. The OML can accomplish CL in more than 200 consecutive classes. It is an online updating method that constantly learns representations and often outperforms rehearsal-based continuous learning methods.
ANML [4]: This is an approach inspired by the neuromodulation process in the brain. This method accomplishes selective activation (forward pass) and selective plasticity (backward pass) in DNNs by meta-learning through a CL process within the prediction network ($f_{\varphi P}$). Optimizing activation timing by controlling the activation of a predictive model based on input conditions enhances continual learning. It is often reported to be superior to OML, as it can be optimized to reduce both forward and backward interference.
@@
OML+AIM [31]: This is a Meta-CL approach that integrates Attentive Independent Mechanisms (AIM), which is a modular component into the OML framework. AIM increases the speed of learning a new task by selecting out the mechanism that best explains the representation of the inner loop to activate.
ANML+AIM [31]: This is a Meta-CL approach that incorporates AIM into the ANML framework. The accuracy of SOTA Meta-CL continues to improve beyond the performance of the previous three approaches by integrating AIM into the existing CL framework.
LifeLearner [26]: This module is based on ANML and combines both lossy and lossless compression to achieve high compression rates and minimize footprint. The LifeLearner approach not only achieves optimal CL efficiency but also significantly reduces energy consumption and latency compared to the SOTA. And it achieves efficient learning even on resource-constrained edge devices.
@@
III-C Datasets
To complement the relative simplicity, limited variety, and lack of generalizability of the datasets tested in previous studies, we added three audio datasets to the single-label image datasets (CIFAR-100 [24], MiniImageNet [52]) from previous studies. Specifically, we used the UrbanSound8K [47] dataset with fewer classes, the ESC-50 [39] dataset with more classes, and the GSCv2 [54] dataset for keyword spotting (KWS).
@@ III-C Datasets
Google Speech Commands V2 (GSCV2) [54] is a massive database of one-second voice clips of 30 short words consisting of a solitary spoken English word or ambient noise divided into 35 classes with 105,829 clips. These clips are derived from a limited number of commands and are spoken by various speakers. Each class has 2,424 input data for training and 314 for testing, with 25 classes allocated for meta-training and the remaining 10 for meta-testing. Then, up to 30 samples are provided for each class during meta-testing.
@@
UrbanSound8K is a comprehensive collection for environment sound classification (ESC) applications comprising 8,732 labeled sounds that are no more than four seconds [47]. This single-labeled dataset is categorized into ten classes from the taxonomy in real-world settings, including children playing sounds, street music, gunshots, etc. Referring to previous studies [50], indicate that four features in each audio clip were extracted and resampled to 22 kHz. Among the ten classes, half of them are allocated for meta-training, while the other half are specifically assigned for meta-testing. Up to 30 samples are given for each class during the meta-testing. Based on the first three seconds of each audio segment, the input to the audio is 128 (number of frames) $\times$ 85 (size of the set of four features).
@@
ESC-50 is a smaller volume collection consisting of 2,000 environment sounds in 5-second segments [39]. This single-labeled dataset is categorized into 50 classes and divided into five groups: animals, natural soundscapes, human sound without speech, domestic sounds and urban noises. Out of the 50 classes. 30 are designated for meta-training and the remaining 20 for meta-testing. Then, up to 30 samples are given for each class during meta-testing. As part of the preprocessing procedure, we resampled the ESC-50 samples to 32kHz and generated an input with dimensions of 157 (number of frames) $\times$ 64 (features) [8, 40].
@@
CIFAR-100 consists of 60,000 images distributed across 100 classes and is widely used for training and assessing various machine learning algorithms in tasks related to image classification [24]. Each class contains 500 photographs for training and 100 images for testing. Of the 100 classes, 70 are designated for meta-training and 30 for meta-testing. In both phases, 30 random training images per class are utilized. For the meta-testing phase, 900 samples are utilized during meta-testing to carry out CL.
@@
MiniImageNet is excerpted from the Imagenet dataset and is utilized in few-shot learning research [52]. Based on the previous research, the MiniImageNet dataset is composed of 64 classes designated for meta-training and 20 for meta-testing. Each class offers 540 images for training and 60 for testing. Moreover, a total of 600 samples are available during the meta-testing phase.
III-D Model Architecture